OpenStudy (anonymous):

What is the value of the expression? –2 + 6 • (–10) A. –62 B. –40 C. 40 D. 120

OpenStudy (anonymous):

A

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@CiaraBelen

OpenStudy (anonymous):

@e.cociuba

OpenStudy (mathstudent55):

Multiply first, then add.

OpenStudy (anonymous):

@BurrDaBoss

OpenStudy (e.cociuba):

-2 + 6 * (-10) = -2 + (- 60) = -2 - 60 = -62

OpenStudy (anonymous):

Well , someone already answered it but , it's A .

OpenStudy (e.cociuba):

ANSWER: A.

OpenStudy (anonymous):

thx

OpenStudy (e.cociuba):

np:)

OpenStudy (mathstudent55):

The order of operations is remembered by this: PEMDAS 1. Parentheses 2. Exponents 3. Multiplications and Divisions from left to right. 4. Additions and Subtractions from left to right. Go down the list, and you see the first operation you need to do is multiplication (item 3.) then addition (item 4.) Do the multiplication first: –2 + 6 • (–10) = -2 + (-60) = -62